OpenStudy (anonymous):

Simplify: 3√5 + √5 Option A: 3√5 Option B: 4√5 Option C: 4√10 Option D: 3√10

terenzreignz (terenzreignz):

Treat sqrt(5) as you would a variable x. What's 3x + x ?

OpenStudy (anonymous):

There I added the options to choose from. And if I go by what your saying, then x =5. 3(5) + 5 = 20. 20 is not an option. Plus, I'm not wanting to solve it. It says to Simplify.

terenzreignz (terenzreignz):

I did not say x = 5 I said x = sqrt (5) You know... square root of 5?

OpenStudy (anonymous):

Right so 3(2.236) + 2.236 = 8.944. So, 8.9? It still doesn't help me. Either you aren't getting the question, or I'm not understanding you. And, again, simplify.

terenzreignz (terenzreignz):

Why put in 2.236? That's not even equal to the square root of 5, exactly. Also, I told you to solve 3x + x That's just 4x. And then replace x with sqrt(x).

terenzreignz (terenzreignz):

*sqrt(5) sorry.

OpenStudy (anonymous):

Right so I'm getting nowhere. Forget it. I'll do research online, hope for somebody else to respond.

terenzreignz (terenzreignz):

/sigh/ \[\huge 3x + x = 4x\]\[\huge 3\sqrt5 + \sqrt5 = 4\sqrt5\]

OpenStudy (anonymous):

Right. Thank you so much, genius. Thank you for sighing and giving me the answer rather than explaining it clearer. Worst experience on this website so far. Gratz'.

terenzreignz (terenzreignz):

I did tell you how to go about it, I said to treat the radical square root of 5 as you would a variable, x. Oh well... Try not to be so rude, next time.

OpenStudy (anonymous):

Try not to be such a smartass next time then. Treating me like I'm an idiot or something instead of explaining it further. Bye.

terenzreignz (terenzreignz):

Good luck with your studies.

OpenStudy (anonymous):

Yeah. Thanks for your 11/10 assistance bud.